A manufacturer produces books at a daily cost of $0.85 per book
and sells them for $1.10 per book. The daily operational overhead is
$400. Determine the break-even point.

Answer by Theo(13342)

profit = revenue minus expense.
profit = 1.10 * x - .85 * x - 400
the break even point is when profit = 0.
equation becomes 0 = 1.10 * x - .85 * x - 400
simplify to get 0 = .25 * x - 400
solve for x to get x = 400 / .25 = 1600
break even is when 1600 books are sold in a day.
you get:
profit = 1.10 * 1600 - .85 * 1600 - 400 = 0
they need to sell 1600 books a day to break even.